# Energy-Efficient Routing Protocol in Vehicular Opportunistic Networks: A Dynamic Cluster-based Routing Using Deep Reinforcement Learning

### v3

@@ I Introduction
Smart city is a concept that aims to integrate information communication with the Internet of Things (IoT) to manage city resources efficiently [A1]. The integration of intelligent vehicles into smart cities can improve safety, traffic flow, and environmental sustainability [A2]. The development of wireless technologies has enabled the implementation of communication systems that rely on vehicles, leading to the emergence of the Vehicular Ad-hoc Network (VANET). VANETs are a type of Mobile Ad Hoc Networks (MANETs) that are differentiated by their highly dynamic network connectivity and frequent topology changes [A3], [A4].
By enabling Intelligent Transport Systems (ITS) to enhance their safety performance and optimize traffic conditions, VANET has established a significant advancement in the field [A5]. This network often leads to a highly dispersed network structure due to the high nodes’ mobility that results in intermittent connectivity [A6]. These networks are known as Opportunistic Networks (OppNets) or Delay Tolerant Networks (DTNs). Note that in the paper, we use the term OppNets rather than DTNs for consistency. A fundamental communication paradigm in OppNets is the Store-Carry-Forward (SCF) architecture (Figure 1). In this model, when a direct communication path between the source and destination is unavailable, data packets are stored locally on a node, carried as the node moves, and then forwarded when the node encounters another that can take the packet closer to its destination. This mechanism either floods the data packets or limit the number of replicas across the network to achieve a high network performance. In the case of flooding, it imposes a high overhead to the network. Moreover, if the buffer of nodes is limited, it increases the packet losses and negatively affecting the delivery ratio. On the other hand, limiting the number of replicas reduces the delivery ratio and increases delays, as data packets have fewer forwarding opportunities.
Cluster-based techniques also address these issues by limiting forwarding opportunities to nodes that (i) have a higher capacity to connect with multiple nodes simultaneously, and (ii) possess sufficient buffer space and energy to carry and forward data packets with lower overhead and delays. However, the existing cluster-based methods often struggle with dynamic environments. They lack mechanisms to adapt to rapidly changing network conditions that leads to inefficient cluster head selection, high overhead, and poor energy management [A9], [A10], [A11].
@@
  • We propose a novel routing protocol that incorporates an AC algorithm to intelligently and optimally select cluster heads in OppNets, based on real-time parameters such as encounter rate, residual energy, and buffer size. Unlike existing clustering techniques [A19], [S1], [S2], which often suffer from workload imbalance due to repeatedly selecting the same nodes as cluster heads, our approach dynamically evaluates node capabilities to ensure balanced resource utilization, improved routing stability, and enhanced network performance in highly dynamic vehicular environments.
  • To address the limitations of fixed-distance clustering in existing methods [New1], [A19], [RN1], [New3], we propose a heuristic function to dynamically adjust the degree of overlap between clusters based on node density and spatial distribution. This will result in improving both scalability and connectivity.
@@ II-A Clustering and Machine Learning-Based Routing in OppNets
In [S1], the authors propose a RL-based Mobility Adaptive Routing protocol for VANETs that optimizes paths using real-time mobility and traffic parameters, achieving improved delivery performance despite implementation complexity challenges. Similarly, in [S2], Kumar et al. introduced RLProph, a RL–based dynamic Programming protocol for opportunistic IoT routing that optimizes delivery probability and latency through adaptive decisions on node connectivity, though its applicability is limited by computational overhead. In [S3], a Latency-Aware Reinforced Routing protocol for OppNets leverages past transmission experiences to adaptively select forwarding paths, thereby reducing E2E delay and improving delivery in dynamic topologies; however, its continuous learning and frequent state updates impose significant computational and energy overhead, limiting suitability for resource-constrained nodes.
@@
In [A19], the authors introduce QUERA, which integrates Q-learning into the Routing Protocol for Low-Power and Lossy Networks (RPL) to enable adaptive parent selection based on energy, link quality, and mobility. This design enhances packet delivery and reduces delay, particularly in mobile IoT environments. The method requires Q-table maintenance, which adds some communication and computation overhead, yet this trade-off highlights an interesting balance between adaptability and resource efficiency that could inspire further optimization. A Reliable and Mobility-Aware RPL (ARMOR) is proposed in [New2] that incorporates a Time-to-Reside (TTR) metric derived from velocity and positional data to better estimate connectivity duration. This mechanism significantly improves packet delivery in low-mobility scenarios. Its evaluation is mainly focused on such environments, leaving opportunities for future exploration of its applicability in higher-mobility conditions, such as vehicular networks, and for strategies to manage the added control signalling more efficiently.
@@
In [New1], the authors propose Energy Efficient Hybrid Clustering and Hierarchical Routing (EEHCHR), which employs Fuzzy C-Means (FCM) clustering based on residual energy and node-to-base station distance to form clusters with both direct and central cluster heads. This design effectively reduces overhead and extends network lifetime. The approach, however, is primarily evaluated under static node deployment, which may be less representative of highly dynamic OppNets scenarios, offering an opportunity for further exploration. In another work [N4], the Energy-efficient technique for selection of Optimum Number of Cluster Head and Grid Head (EOCGS) protocol proposes a hybrid clustering approach that also leverages FCM, residual energy, and base station distance to selectively form clusters, thereby reducing overhead. While effective, the protocol’s reliance on predefined grid structures means its adaptability to changing or mobile environments could be further investigated. A modified version of Residual Energy based Cluster Head Selection (MRCH) in Low Energy Adaptive Clustering Hierarchy (RCH-LEACH) is proposed in [N5], which improves energy efficiency by selecting cluster heads based on node energy levels and optimizing the cluster head percentage. MRCH extends network lifetime and shows robustness under different base station placements. Its cluster-rotating mechanism is well suited for stable deployments, while future work could consider enhancements for scenarios with higher mobility and rapidly evolving topologies. A Cluster-based Routing using Meta-Heuristic Algorithms for VANET (CRMHA-VANET) is proposed in [New3], which combines Gray Wolf Optimization (GWO) for cluster head selection and Technique for Order of Preference by Similarity to Ideal Solution (TOPSIS) for route selection, optimizing intra-cluster distance, link quality, and load balancing. Although it achieves good performance, its dependence on RSUs limits scalability in OppNets with sparse infrastructure. In [RN3], a cluster-based routing scheme employing Harris Hawks Optimization (HHO) is proposed, where cluster head selection is guided by node mobility, residual energy, and transmission distance. This approach enhances the delivery ratio and reduces latency. HHO-based clustering offers clear benefits for cluster-head selection and load balancing, though factors such as computational overhead and adaptability in dynamic environments remain areas for further consideration.
@@
Moghaddasi et al. [RN1] propose a Double Deep Q-Network (DDQN) for 5G-enabled Vehicular Edge Computing (VEC), which reduces energy consumption and latency through adaptive task offloading. However, its reliance on stable edge infrastructure suggests opportunities for further evaluation in highly dynamic environments such as OppNets. In [RN2], a DRL-based task offloading method is proposed for three-layer Device-to-Device (D2D)–edge–cloud IoT architecture, employing advanced neural networks to optimize offloading decisions across heterogeneous environments. The approach achieves significant improvements in energy efficiency, latency reduction, and utility enhancement through adaptive real-time decision-making. However, future research could explore lightweight implementation strategies to optimize computational requirements for broader deployment across diverse IoT infrastructures. In [A41], the authors propose ECRDP, an Efficient Clustering Routing approach that integrates Density Peaks Clustering (DPC) with Particle Swarm Optimization (PSO) to adapt routing decisions according to vehicle density, traffic conditions, road topology, and link quality. While the method demonstrates adaptability, its high computational complexity results in suboptimal delivery ratios and throughput.
@@
Existing protocols face challenges in cluster head selection, delivery ratio, E2E delay, and energy efficiency, particularly in highly dynamic OppNets. RL-based methods [S1], [S2], [S3], [A19], and [New2] offer adaptability but incur computational overhead, while clustering approaches [New1], [New3], [RN3], [N4], [N5], [A41], and [RN1] struggle with static thresholds or infrastructure dependency. CR-DRL addresses these gaps by integrating an AC framework for adaptive cluster head selection and a heuristic function for dynamic cluster overlap, enhancing scalability, energy efficiency, and compatibility with TCP/IP protocols without heavy infrastructure reliance. Table I provides a detailed comparison of these methods’ strengths, limitations, and metrics.
@@
[A19] Q-learning Adapts to mobility; improves energy efficiency, packet delivery, and reduced delay. Requires Q-table maintenance; increased memory and processing demand. Packet delivery ratio, delay.
[S1] RL Dynamic path learning; better delivery in VANETs. High training overhead; complex reward design. Delivery ratio, latency.
[S2] RL Effective in intermittent links; improves delivery. Computationally intensive; higher overhead. Delivery rate, latency.
[New1] FCM Balanced energy use; lower overhead; extended lifetime. Static thresholds; lacks adaptability. Network lifetime, energy consumption.
[RN1] DDQN Adaptive offloading; reduced energy and latency. Assumes stable edge infrastructure; unsuitable for OppNets. Energy consumption, latency.
[New3] GWO Multi-objective fitness function; hierarchical structure for stability; balanced clustering. Increased packet overhead; RSU dependency; scalability concerns. Delivery rate; throughput; E2E delay.
[RN3] HHO Improved delivery ratio; reduced delay. Unstable convergence; needs parameter tuning. Delivery ratio, delay.
[N4] FCM Balanced energy use; prolonged network life. Grid structure limits adaptability. Energy usage, network lifetime
[N5] Clustering Lifetime extension; robust under various BS locations. Static rotation; not suited for mobile/dynamic cases. Energy efficiency, network lifetime.
[A41] DPC + PSO Context-aware routing; handles density/topology variations. High complexity; underperforms in delivery ratio. Delivery ratio, throughput.
[S3] RL Reduces E2E delay; adapts to topology changes. High learning cost; not ideal for constrained devices. E2E delay, delivery rate.
[New2] TTR Improved delivery ratio; stable connection establishment; comparable power consumption. Control overhead increase; limited to low-speed mobility; potential latency increase. Delivery rate; power consumption, E2E delay.
[RN2] DRL Improved energy efficiency; latency reduction and utility enhancement. Complexity; high learning cost Delivery ratio, delay, energy efficiency.
@@ III-A Preliminary
In this section, we present the RL strategy adopted for routing in OppNets, focusing on the AC framework [AC1], an advanced variant of the policy gradient with a baseline algorithm. This framework effectively tackles the challenges of dynamic environments like OppNets, where traditional routing techniques often fail due to high node mobility and intermittent connectivity. The AC framework is formulated as a Markov Decision Process (MDP), represented by the tuple ($s,a,P,R,\gamma$) [N1], where:
@@
Where $J(\theta)$ is the performance objective, ${\pi}_{\theta}(a|s)$ represents the policy parameterized by $\theta$, which defines the probability of taking action $a$ in state $s$ and $R_{t}$ is the reward received at time step $t$. [A50]. The policy gradient theorem [N2] provides the gradient of the objective function:
@@
Where $Q^{\pi}(s,a)$ is the action-value function, representing the expected return after taking action $a$ in state $s$ under policy $\pi$. While the above gradient works, it often suffers from high variance, especially in dynamic environments such as OppNets. To mitigate this issue, a baseline function $b(s)$, typically the state-value function $V(s;\omega)$, is proposed [A51]:
@@
Where $Q^{\pi}(s,a)-V(s;\omega)$ is the advantage function, measuring how much better an action is compared to the baseline [A51]:
@@
Where $\alpha_{a}$ is the actor’s learning rate, and $A(s,a)$ guides the update based on the advantage of the chosen action. The critic updates the value function parameters $\omega$ using the Temporal-Difference (TD) error $\delta_{t}$ [A38]:
@@ III-B System description
In our network, every node utilizes a Multiple Input Multiple Output (MIMO) channel to enhance communication quality among vehicles. When a vehicle is a common member of multiple clusters, it can establish communication with other cluster heads. To ensure higher transmission speed and reliable communication between vehicles, we have implemented MIMO $4$ in this system [N3]. Lastly, clusters should have minimal communication with other clusters to facilitate communication through shared members. In Table II, we have provided a summary of the notations used in the CR-DRL method.
@@ III-C Problem formulation
$\tau_{t}$ is the reliability of the link $G_{i}$} at the time t{t}. ${T}_{e}$ is estimated availability time, and f(T){f}(T) is the probability density function of communication [A28] duration T{T} of ${v}_{i}$ and ${v}_{j}$. Traditional routing strategies struggle to ensure reliable data transmission in such an environment, making it crucial to develop an adaptive approach that maximizes delivery efficiency while minimizing energy consumption and transmission delays. To address this challenge, we frame the routing problem as an optimization task that seeks to enhance network performance while maintaining resource efficiency. To optimize data transmission in vehicular networks, we maximize delivery ratio and throughput, which ensure efficient data forwarding, while minimizing E2E delay to reduce transmission latency. The optimization problem is formulated as follows:

### v4

@@ I Introduction
Smart city is a concept that aims to integrate information communication with the Internet of Things (IoT) to manage city resources efficiently [32]. The integration of intelligent vehicles into smart cities can improve safety, traffic flow, and environmental sustainability [35]. The development of wireless technologies has enabled the implementation of communication systems that rely on vehicles, leading to the emergence of the Vehicular Ad-hoc Network (VANET). VANETs are a type of Mobile Ad Hoc Networks (MANETs) that are differentiated by their highly dynamic network connectivity and frequent topology changes [39], [2].
By enabling Intelligent Transport Systems (ITS) to enhance their safety performance and optimize traffic conditions, VANET has established a significant advancement in the field [31]. This network often leads to a highly dispersed network structure due to the high nodes’ mobility that results in intermittent connectivity [28]. These networks are known as Opportunistic Networks (OppNets) or Delay Tolerant Networks (DTNs). Note that in the paper, we use the term OppNets rather than DTNs for consistency. A fundamental communication paradigm in OppNets is the Store-Carry-Forward (SCF) architecture (Figure 1). In this model, when a direct communication path between the source and destination is unavailable, data packets are stored locally on a node, carried as the node moves, and then forwarded when the node encounters another that can take the packet closer to its destination. This mechanism either floods the data packets or limit the number of replicas across the network to achieve a high network performance. In the case of flooding, it imposes a high overhead to the network. Moreover, if the buffer of nodes is limited, it increases the packet losses and negatively affecting the delivery ratio. On the other hand, limiting the number of replicas reduces the delivery ratio and increases delays, as data packets have fewer forwarding opportunities.
Cluster-based techniques also address these issues by limiting forwarding opportunities to nodes that (i) have a higher capacity to connect with multiple nodes simultaneously, and (ii) possess sufficient buffer space and energy to carry and forward data packets with lower overhead and delays. However, the existing cluster-based methods often struggle with dynamic environments. They lack mechanisms to adapt to rapidly changing network conditions that leads to inefficient cluster head selection, high overhead, and poor energy management [22], [1], [29].
@@
  • We propose a novel routing protocol that incorporates an AC algorithm to intelligently and optimally select cluster heads in OppNets, based on real-time parameters such as encounter rate, residual energy, and buffer size. Unlike existing clustering techniques [30], [37], [34], which often suffer from workload imbalance due to repeatedly selecting the same nodes as cluster heads, our approach dynamically evaluates node capabilities to ensure balanced resource utilization, improved routing stability, and enhanced network performance in highly dynamic vehicular environments.
  • To address the limitations of fixed-distance clustering in existing methods [23], [30], [19], [15], we propose a heuristic function to dynamically adjust the degree of overlap between clusters based on node density and spatial distribution. This will result in improving both scalability and connectivity.
@@ II-A Clustering and Machine Learning-Based Routing in OppNets
In [37], the authors propose a RL-based Mobility Adaptive Routing protocol for VANETs that optimizes paths using real-time mobility and traffic parameters, achieving improved delivery performance despite implementation complexity challenges. Similarly, in [34], Kumar et al. introduced RLProph, a RL–based dynamic Programming protocol for opportunistic IoT routing that optimizes delivery probability and latency through adaptive decisions on node connectivity, though its applicability is limited by computational overhead. In [33], a Latency-Aware Reinforced Routing protocol for OppNets leverages past transmission experiences to adaptively select forwarding paths, thereby reducing E2E delay and improving delivery in dynamic topologies; however, its continuous learning and frequent state updates impose significant computational and energy overhead, limiting suitability for resource-constrained nodes.
@@
In [30], the authors introduce QUERA, which integrates Q-learning into the Routing Protocol for Low-Power and Lossy Networks (RPL) to enable adaptive parent selection based on energy, link quality, and mobility. This design enhances packet delivery and reduces delay, particularly in mobile IoT environments. The method requires Q-table maintenance, which adds some communication and computation overhead, yet this trade-off highlights an interesting balance between adaptability and resource efficiency that could inspire further optimization. A Reliable and Mobility-Aware RPL (ARMOR) is proposed in [20] that incorporates a Time-to-Reside (TTR) metric derived from velocity and positional data to better estimate connectivity duration. This mechanism significantly improves packet delivery in low-mobility scenarios. Its evaluation is mainly focused on such environments, leaving opportunities for future exploration of its applicability in higher-mobility conditions, such as vehicular networks, and for strategies to manage the added control signalling more efficiently.
@@
In [23], the authors propose Energy Efficient Hybrid Clustering and Hierarchical Routing (EEHCHR), which employs Fuzzy C-Means (FCM) clustering based on residual energy and node-to-base station distance to form clusters with both direct and central cluster heads. This design effectively reduces overhead and extends network lifetime. The approach, however, is primarily evaluated under static node deployment, which may be less representative of highly dynamic OppNets scenarios, offering an opportunity for further exploration. In another work [24], the Energy-efficient technique for selection of Optimum Number of Cluster Head and Grid Head (EOCGS) protocol proposes a hybrid clustering approach that also leverages FCM, residual energy, and base station distance to selectively form clusters, thereby reducing overhead. While effective, the protocol’s reliance on predefined grid structures means its adaptability to changing or mobile environments could be further investigated. A modified version of Residual Energy based Cluster Head Selection (MRCH) in Low Energy Adaptive Clustering Hierarchy (RCH-LEACH) is proposed in [36], which improves energy efficiency by selecting cluster heads based on node energy levels and optimizing the cluster head percentage. MRCH extends network lifetime and shows robustness under different base station placements. Its cluster-rotating mechanism is well suited for stable deployments, while future work could consider enhancements for scenarios with higher mobility and rapidly evolving topologies. A Cluster-based Routing using Meta-Heuristic Algorithms for VANET (CRMHA-VANET) is proposed in [15], which combines Gray Wolf Optimization (GWO) for cluster head selection and Technique for Order of Preference by Similarity to Ideal Solution (TOPSIS) for route selection, optimizing intra-cluster distance, link quality, and load balancing. Although it achieves good performance, its dependence on RSUs limits scalability in OppNets with sparse infrastructure. In [14], a cluster-based routing scheme employing Harris Hawks Optimization (HHO) is proposed, where cluster head selection is guided by node mobility, residual energy, and transmission distance. This approach enhances the delivery ratio and reduces latency. HHO-based clustering offers clear benefits for cluster-head selection and load balancing, though factors such as computational overhead and adaptability in dynamic environments remain areas for further consideration.
@@
Moghaddasi et al. [19] propose a Double Deep Q-Network (DDQN) for 5G-enabled Vehicular Edge Computing (VEC), which reduces energy consumption and latency through adaptive task offloading. However, its reliance on stable edge infrastructure suggests opportunities for further evaluation in highly dynamic environments such as OppNets. In [18], a DRL-based task offloading method is proposed for three-layer Device-to-Device (D2D)–edge–cloud IoT architecture, employing advanced neural networks to optimize offloading decisions across heterogeneous environments. The approach achieves significant improvements in energy efficiency, latency reduction, and utility enhancement through adaptive real-time decision-making. However, future research could explore lightweight implementation strategies to optimize computational requirements for broader deployment across diverse IoT infrastructures. In [12], the authors propose ECRDP, an Efficient Clustering Routing approach that integrates Density Peaks Clustering (DPC) with Particle Swarm Optimization (PSO) to adapt routing decisions according to vehicle density, traffic conditions, road topology, and link quality. While the method demonstrates adaptability, its high computational complexity results in suboptimal delivery ratios and throughput.
@@
Existing protocols face challenges in cluster head selection, delivery ratio, E2E delay, and energy efficiency, particularly in highly dynamic OppNets. RL-based methods [37], [34], [33], [30], and [20] offer adaptability but incur computational overhead, while clustering approaches [23], [15], [14], [24], [36], [12], and [19] struggle with static thresholds or infrastructure dependency. CR-DRL addresses these gaps by integrating an AC framework for adaptive cluster head selection and a heuristic function for dynamic cluster overlap, enhancing scalability, energy efficiency, and compatibility with TCP/IP protocols without heavy infrastructure reliance. Table I provides a detailed comparison of these methods’ strengths, limitations, and metrics.
@@
[30] Q-learning Adapts to mobility; improves energy efficiency, packet delivery, and reduced delay. Requires Q-table maintenance; increased memory and processing demand. Packet delivery ratio, delay.
[37] RL Dynamic path learning; better delivery in VANETs. High training overhead; complex reward design. Delivery ratio, latency.
[34] RL Effective in intermittent links; improves delivery. Computationally intensive; higher overhead. Delivery rate, latency.
[23] FCM Balanced energy use; lower overhead; extended lifetime. Static thresholds; lacks adaptability. Network lifetime, energy consumption.
[19] DDQN Adaptive offloading; reduced energy and latency. Assumes stable edge infrastructure; unsuitable for OppNets. Energy consumption, latency.
[15] GWO Multi-objective fitness function; hierarchical structure for stability; balanced clustering. Increased packet overhead; RSU dependency; scalability concerns. Delivery rate; throughput; E2E delay.
[14] HHO Improved delivery ratio; reduced delay. Unstable convergence; needs parameter tuning. Delivery ratio, delay.
[24] FCM Balanced energy use; prolonged network life. Grid structure limits adaptability. Energy usage, network lifetime
[36] Clustering Lifetime extension; robust under various BS locations. Static rotation; not suited for mobile/dynamic cases. Energy efficiency, network lifetime.
[12] DPC + PSO Context-aware routing; handles density/topology variations. High complexity; underperforms in delivery ratio. Delivery ratio, throughput.
[33] RL Reduces E2E delay; adapts to topology changes. High learning cost; not ideal for constrained devices. E2E delay, delivery rate.
[20] TTR Improved delivery ratio; stable connection establishment; comparable power consumption. Control overhead increase; limited to low-speed mobility; potential latency increase. Delivery rate; power consumption, E2E delay.
[18] DRL Improved energy efficiency; latency reduction and utility enhancement. Complexity; high learning cost Delivery ratio, delay, energy efficiency.
@@ III-A Preliminary
In this section, we present the RL strategy adopted for routing in OppNets, focusing on the AC framework [5], an advanced variant of the policy gradient with a baseline algorithm. This framework effectively tackles the challenges of dynamic environments like OppNets, where traditional routing techniques often fail due to high node mobility and intermittent connectivity. The AC framework is formulated as a Markov Decision Process (MDP), represented by the tuple ($s,a,P,R,\gamma$) [11], where:
@@
Where $J(\theta)$ is the performance objective, ${\pi}_{\theta}(a|s)$ represents the policy parameterized by $\theta$, which defines the probability of taking action $a$ in state $s$ and $R_{t}$ is the reward received at time step $t$. [26]. The policy gradient theorem [6] provides the gradient of the objective function:
@@
Where $Q^{\pi}(s,a)$ is the action-value function, representing the expected return after taking action $a$ in state $s$ under policy $\pi$. While the above gradient works, it often suffers from high variance, especially in dynamic environments such as OppNets. To mitigate this issue, a baseline function $b(s)$, typically the state-value function $V(s;\omega)$, is proposed [38]:
@@
Where $Q^{\pi}(s,a)-V(s;\omega)$ is the advantage function, measuring how much better an action is compared to the baseline [38]:
@@
Where $\alpha_{a}$ is the actor’s learning rate, and $A(s,a)$ guides the update based on the advantage of the chosen action. The critic updates the value function parameters $\omega$ using the Temporal-Difference (TD) error $\delta_{t}$ [16]:
@@ III-B System description
In our network, every node utilizes a Multiple Input Multiple Output (MIMO) channel to enhance communication quality among vehicles. When a vehicle is a common member of multiple clusters, it can establish communication with other cluster heads. To ensure higher transmission speed and reliable communication between vehicles, we have implemented MIMO $4$ in this system [10]. Lastly, clusters should have minimal communication with other clusters to facilitate communication through shared members. In Table II, we have provided a summary of the notations used in the CR-DRL method.
@@ III-C Problem formulation
$\tau_{t}$ is the reliability of the link $G_{i}$} at the time t{t}. ${T}_{e}$ is estimated availability time, and f(T){f}(T) is the probability density function of communication [13] duration T{T} of ${v}_{i}$ and ${v}_{j}$. Traditional routing strategies struggle to ensure reliable data transmission in such an environment, making it crucial to develop an adaptive approach that maximizes delivery efficiency while minimizing energy consumption and transmission delays. To address this challenge, we frame the routing problem as an optimization task that seeks to enhance network performance while maintaining resource efficiency. To optimize data transmission in vehicular networks, we maximize delivery ratio and throughput, which ensure efficient data forwarding, while minimizing E2E delay to reduce transmission latency. The optimization problem is formulated as follows: